OpenStudy (anonymous):

Could someone explain to me why is it important to simplify radical expressions before adding or subtracting?

OpenStudy (across):

It is not really important, but you can say that it makes computation easier.

OpenStudy (anonymous):

yes, what @across said. for example, sqrt8 + sqrt50. if you needed this value, would you rather work with (sqrt8 +sqrt50) or just 7sqrt2?

OpenStudy (anonymous):

Consider this example Add: \[\sqrt{s} \sqrt{?}\] \[\sqrt{s} \sqrt{?}\] Start with the given expression. Can you add these two radicals in this state? The answer is simply no. We need to simplify the radicals. Take the square root of 4 to get 2. Take the square root of 16 to get 4. In this step we have simplified the radicals. Add 2+4 So simplifies to 6. In other words,